Please help me solve this:
Suppose $1000 is invested at an annual rate of 5.7% compounded continuously.
A.What is the ammount after 6 years?
B.How long would it take for the ammount in the account to quadruple?

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(48555)

Suppose $1000 is invested at an annual rate of 5.7% compounded continuously.
A.What is the ammount after 6 years?
A(t) = Pe^(rt)
A(6) = 1000e^(0.057*6)
A(6) = 1000*1.4078
A(6) = $1407.80
-------------------------
B.How long would it take for the ammount in the account to quadruple?
4000 = 1000e^(0.057t)
4 = e^(0.057t)
Take the natural log of both sides.
0.057t = ln4
t = [ln4/0.057]
t = 24.32 years
================================================
Cheers,
Stan H.

Answer by scott8148(5891)

A=Pe^(rt)

A=1000e^(.057*6)


let t="time to quadruple"
4=e(.057*t) __ taking natural log __ ln(4)=.057*t __ dividing by .057 __ (ln(4))/.057=t